The average age of 5 children is 8 years. If the age of the father be included, the average is
increased by 7 years. Find the age of the father.

Answer by htmentor(1343)

The average age of 5 children is 8 years. If the age of the father be included, the average is
increased by 7 years. Find the age of the father.
Let S = the sum of the children's ages
The average of the 5 children's ages = 8 = S/5 -> S = 40
Let the father's age = f
The average, including the father = 8 + 7 = 15 = (S+f)/6
Since S=40, we have
15 = (40+f)/6
f = 90 - 40 = 50
So the father is 50 years old